OpenStudy (anonymous):

The volume of a cube is 36 cubic feet. What is the volume of a pyramid that fits exactly inside the cube? Provide an explanation and proof for your answer

OpenStudy (nirmalnema):

volume of cube =a^3 side of cube = \[\sqrt[3]{36}\]

OpenStudy (nirmalnema):

which is equal to 3.3

OpenStudy (anonymous):

is that everything I right? lol jw

OpenStudy (nirmalnema):

no it wasn't/

OpenStudy (anonymous):

wgat else do i write?